# SlumpGuard: An AI-Powered Real-Time System for Automated Concrete Slump Prediction via Video Analysis

### v2

@@ 4.3.3 Model
We employed the ResNet 3D architecture [26] as our video classification backbone, designed to jointly capture spatial and temporal patterns. We experimented with two variants: Mixed Convolution and (2+1)D convolution. The Mixed Convolution model uses 3D convolutions in early layers for motion modeling and 2D convolutions in later layers for efficient spatial reasoning, achieving performance comparable to full 3D models with fewer parameters. The (2+1)D model factorizes 3D convolutions into separate spatial and temporal steps, enhancing optimization and model capacity.
@@ 4.3.4 Training
To formulate the slump estimation task as a classification problem, we categorized the continuous slump values into five categorical intervals (e.g. 150–180mm, 180–210mm), where the width of each interval reflects the $\pm 30$ mm tolerance specified in the KS F 4009 standard [38], along with an additional margin for human error. For intervals at the extremes (i.e. below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
@@
where $y_{i}^{soft}\in[0,1]$ denotes the soft label vector obtained via MixUp and label smoothing, and $\hat{y}_{i}$ represents the predicted probability for class $i$.
@@ 5.1 Dataset
Over an eight-month period, a dataset was constructed by replicating real-world construction site conditions in a controlled experimental settings. The recorded videos were segmented into 10-second intervals, resulting in $4,504$ training samples, $998$ validation samples, and $941$ test samples, as shown in Figure 4.
@@ 5.4.1 Model Training Setup
We trained the model using 16-frame video clips, sampled at a frame interval of 2, and extended by a temporal factor of 2. The resize ratio was randomly selected from { 3/4, 4/3 }, and standard data augmentations were applied. For spatial augmentation, we used random horizontal flipping with a probability of 0.5 and ColorJitter with the following ranges: brightness 0.4, contrast 0.4, saturation 0.4, and hue 0.1. Additionally, MixUp ($\alpha$ = 0.2, probability = 0.5) and CutMix ($\alpha$ = 0.2, probability = 0.5) were employed to generate soft-label training samples, encouraging better generalization and robustness to noisy data. Label smoothing with a factor, $\lambda_{cls}$, of 0.1 was also applied.
The model was optimized using the AdamW optimizer [39] with a learning rate of 1e41e-4, and a weight decay of 1e41e-4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [40] with cosine annealing, with a peak learning rate set to 1e31e-3.
@@
Transformer TimeSFormer [28] 0.6980 0.7280 0.5724 0.6155
@@ 5.4.2 Results
As shown in Table 3, our models employing ResNet-3D as the backbone achieve over 80% accuracy on both the validation and test sets. In contrast, the Transformer [41]-based TimeSFormer [28] exhibits lower accuracy. Among the evaluated architectures, we selected R(2+1)D-18 as our final model, as it achieved the highest performance on both validation and test sets. Based on this architecture, we conducted a detailed ablation study, and the corresponding results are reported in Table 3.

### v3

@@ 4.3.3 Model
We employed the ResNet 3D architecture [tran2018:closer] as our video classification backbone, designed to jointly capture spatial and temporal patterns. We experimented with two variants: Mixed Convolution and (2+1)D convolution. The Mixed Convolution model uses 3D convolutions in early layers for motion modeling and 2D convolutions in later layers for efficient spatial reasoning, achieving performance comparable to full 3D models with fewer parameters. The (2+1)D model factorizes 3D convolutions into separate spatial and temporal steps, enhancing optimization and model capacity.
@@ 4.3.4 Training
To formulate the slump estimation task as a classification problem, we categorized the continuous slump values into five categorical intervals (e.g. 150–180mm, 180–210mm), where the width of each interval reflects the $\pm 30$ mm tolerance specified in the KS F 4009 standard [ksf4009], along with an additional margin for human error. For intervals at the extremes (i.e. below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
@@
where $y_{i}^{soft}\in[0,1]$ denotes the soft label vector obtained via MixUp and label smoothing, and $\hat{y}_{i}$ represents the predicted probability for class $i$. We empirically set $\lambda_{cls}=0.1$.
@@ 5.1 Dataset
Over an eight-month period, a dataset was constructed by replicating real-world construction site conditions in a controlled experimental settings. The recorded videos were segmented into uniform 10-second intervals, discarding any residual clips shorter than 10 seconds. This process resulted in $4,504$ training samples, $998$ validation samples, and $941$ test samples, as shown in Figure 4.
@@ 5.4.1 Model Training Setup
We trained the model using 16-frame video clips, sampled at a frame interval of 2, and extended by a temporal factor of 2. The resize ratio was randomly selected from { 3/4, 4/3 }, and standard data augmentations were applied. For spatial augmentation, we used random horizontal flipping with a probability of 0.5 and ColorJitter with the following ranges: brightness 0.4, contrast 0.4, saturation 0.4, and hue 0.1. Additionally, MixUp ($\alpha$ = 0.2, probability = 0.5) and CutMix ($\alpha$ = 0.2, probability = 0.5) were employed to generate soft-label training samples, encouraging better generalization and robustness to noisy data. Label smoothing with a factor of 0.1 was also applied.
The model was optimized using the AdamW optimizer [loshchilov2017:decoupled] with a learning rate of 1e41e-4, and a weight decay of 1e41e-4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [smith2019:super] with cosine annealing, with a peak learning rate set to 1e31e-3.
@@
Transformer TimeSFormer [bertasius2021:space] 0.6980 0.7280 0.5724 0.6155
@@ 5.4.2 Results
As shown in Table 3, our models employing ResNet-3D as the backbone achieve over 80% accuracy on both the validation and test sets. In contrast, the Transformer [vaswani2017:attention]-based TimeSFormer [bertasius2021:space] exhibits lower accuracy. Among the evaluated architectures, we selected R(2+1)D-18 as our final model, as it achieved the highest performance on both validation and test sets. Based on this architecture, we conducted a detailed ablation study, and the corresponding results are reported in Table 3.
@@
Figure 6: Screenshot of the expert evaluation interface used in our study. A Google Forms–based questionnaire was provided to domain experts, requiring them to choose a single slump interval from five predefined categories.
5.4.3 Expert Evaluation
To further validate the practical reliability of SlumpGuard under real-world operational conditions, we conducted an expert human evaluation to assess how accurately people can visually estimate concrete slump. For this evaluation, we asked 11 domain experts with over 15 years of field experience to review ten video samples of concrete discharging from mixer trucks and estimate the corresponding slump range based solely on visual observation. All selected videos were cases for which SlumpGuard produced correct predictions, and each clip shown to the experts consisted of a 10-second segment. A screenshot of the Google Forms interface used for this expert evaluation is shown in Figure 6, with the results summarized in Tables LABEL:tab:expert_acc and Figure 7.
Despite being highly experienced professionals, the experts exhibited notable disagreement when visually estimating slump ranges. As reported in Table LABEL:tab:expert_acc, only 9.1–-45.5% of the experts correctly identified the ground-truth slump in each sample, and no video exhibited unanimous agreement among all evaluators. Moreover, the confusion matrix in Figure 7 indicates that most misclassifications occurred between adjacent slump ranges (e.g., 180–210 mm vs. 210–240 mm), revealing the inherent difficulty of visually distinguishing small differences in concrete workability. These findings highlight the subjective variability and inconsistency of human judgment in real-site conditions, even among experts with extensive field experience. In contrast, SlumpGuard provided correct predictions for all evaluated samples, demonstrating its potential as a reliable and objective tool for real-time slump estimation without the need for manual testing.
Table 4: Expert evaluation for slump estimation. Ground-truth slump values and the number (percentage) of correct responses among eleven experts.
GT ($m$) Correct (11)
Under 150 (150) 4 (36.4%)
180 $\sim$ 210 (200) 3 (27.3%)
Under 150 (145) 5 (45.5%)
180 $\sim$ 210 (190) 5 (45.5%)
Over 240 (280) 5 (45.5%)
210 $\sim$ 240 (280) 1 (9.1%)
150 $\sim$ 180 (170) 3 (27.3%)
180 $\sim$ 210 (190) 3 (27.3%)
150 $\sim$ 180 (165) 4 (36.4%)
210 $\sim$ 240 (220) 3 (27.3%)
[Uncaptioned image] Figure 7: Confusion matrix of expert predictions. Numbers represent counts and percentages relative to the ground-truth slump class.
6 Real-World Deployment
Figure 8: Screenshot of the deployed SlumpGuard user interface. The left panel visualizes chute detection, the passing-line based flow gating, and cumulative flow at the chute center. The right panel displays the system’s slump prediction (updated every NN frames), the ordered slump value, and the final majority-voted slump result. A warning message is triggered when the predicted range does not match the ordered specification.
To demonstrate the practical applicability of SlumpGuard in real construction environments, we present the on-site operational user interface (UI) used during deployment, as shown in Figure 8.
Before concrete placement begins, the ordered slump specification for each batch is entered into the system (see “Input” in Figure 8). When two mixer trucks arrive at the placement area, the system automatically detects both chutes without any manual intervention. It then identifies which of the two trucks is actively discharging concrete and determines the precise onset of pouring (see red circle and passing line in Figure 8). Once the pouring onset is detected, the slump classification model is activated and begins predicting the slump category by every $N$ frames (see “Current Slump Value” in Figure 8). The system produces predictions over ten successive inference cycles (i.e., $T=10$), after which the final slump value is determined through majority voting. Based on this prediction, the system compares the inferred slump range with the ordered specification for the corresponding batch. If the prediction deviates from the specified range, the UI displays a warning flag; otherwise, the batch is marked as acceptable.
This interface is lightweight, requires no specialized hardware beyond a fixed camera, and integrates seamlessly into existing construction workflows. This deployed interface illustrates how SlumpGuard functions as a fully autonomous, practical tool for continuous and non-intrusive quality assurance on site, offering an interpretable and actionable output that supports real-time decision-making during concrete placement.
7 Discussion
This section discusses four key aspects of the proposed system: (1) the architectural design of the modular pipeline, (2) the system-level implications for real-world deployment, (3) the practical constraints imposed by actual construction workflows, and (4) the expert-level evaluation that contextualizes the model’s performance relative to human judgment.
7.1 Pipeline Design
The three-stage structure adopted in this work provides several advantages for field deployment. Separating the tasks of chute localization, pouring-event detection, and slump classification allows each component to operate under well-defined assumptions, improving interpretability and enabling targeted refinement when system behavior must be analyzed or adapted to a new site. The motion-gated mechanism ensures that only slump-relevant temporal windows are processed, reducing computational cost and supporting real-time operation. Although an end-to-end formulation may appear conceptually simpler, it is structurally unsuitable for this task. An end-to-end model would need to implicitly learn the pouring onset, accurately localize the chute under large geometric and rotational variation, and simultaneously infer motion cues required for flow-based gating. These sub-tasks demand fundamentally different spatial and temporal inductive biases, which cannot be reliably integrated into a single unified network. By contrast, the modular pipeline offers greater flexibility, maintainability, and transparency, which are essential properties for real-world operation in dynamic on-site environments.
7.2 System-level Implication
The purpose of SlumpGuard extends beyond its modular pipeline design to its ability to function as a fully autonomous system that aligns with practical construction-site workflows. Sensor-based approaches often encounter operational challenges, such as restrictions on hardware installation, differences in responsibility between concrete suppliers and contractors, and additional labor requirements for dedicated monitoring. In contrast, the proposed system requires only a single fixed camera and does not depend on any mixer-truck-mounted equipment, allowing it to be integrated seamlessly into standard concrete placement procedures. This vision-based, non-intrusive design enables deployment across a wide range of site conditions while offering scalability and practical utility for continuous quality monitoring.
Furthermore, explicit modeling of pouring onset and the interpretability of intermediate outputs facilitate inspection of system behavior under various environmental conditions and support the identification of potential error sources, thereby enhancing operational trust. These characteristics demonstrate that SlumpGuard is a field-ready solution that offers the reliability, transparency, and maintainability required for real-world construction environments. They also illustrate how an integrated vision-based pipeline can serve as a practical alternative or complement to conventional manual or sensor-based inspection processes.
7.3 Practical Constraints in Real Construction Workflows
The proposed system reflects the practical constraints under which slump inspection is conducted in real construction environments. In standard concrete placement, mixer trucks are required to remain stationary on level ground, and discharging from a tilted vehicle is not permitted due to safety regulations and the risk of altering fresh concrete properties. Therefore, ground-slope variation does not represent a realistic condition for slump evaluation in practice.
In addition, slump inspection is not carried out at night or under unstable illumination, as concrete placement and visual quality assessment are mandated to occur under sufficiently lit conditions. Since SlumpGuard operates within these same constraints and focuses analysis on the interior of the chute region, unrealistic scenarios such as steep slopes or nighttime pouring fall outside the practical scope of slump evaluation.
7.4 Expert-level Evaluation
The expert study provides additional insight into how visual slump assessment is performed in practice and helps contextualize the model’s performance relative to human judgment. Notably, the variability in expert responses indicates that visual estimation of slump is inherently subjective. In real construction workflows, slump is rarely determined by closely observing the concrete as it flows out of the mixer truck. Instead, field engineers typically inspect the concrete after it has accumulated on the placement surface, relying on experiential cues developed through repeated exposure to that specific condition. As a result, the experts in our study were asked to evaluate a visual scenario that differs from their usual inspection context, namely, assessing slump directly from the discharge stream rather than from the already deposited material.
This mismatch between familiar field practice and the evaluation setting likely contributed to the lower-than-expected agreement among experts. Because their judgments depend heavily on tacit knowledge and accumulated impression-based cues, the absence of familiar reference conditions can lead to divergence in interpretation. The expert-level evaluation thus reflects the inherent ambiguity associated with purely visual slump estimation and underscores the practical challenge of achieving consistent human judgment in such tasks.
Taken together, these findings suggest that the predictive consistency offered by SlumpGuard may provide meaningful support in situations where visual assessment alone leads to substantial inter-operator variability. This further highlights the potential role of automated, vision-based systems in complementing traditional quality assurance workflows.
8 Limitations
This study has several limitations. First, the optical-flow–based motion detection is sensitive to shadows and illumination changes, particularly within darker or low-contrast regions of the chute interior, where flow vectors may become distorted. However, this effect is largely limited to the frame selection and pouring-event detection stages rather than directly impacting the accuracy of slump classification. Furthermore, the use of data augmentation techniques during training contributes to improving the illumination robustness of the final classification model. Second, although the dataset was collected in a controlled experimental environment designed to closely replicate real concrete pouring operations, it does not originate from fully uncontrolled and actively operating construction sites. Additional validation across a broader spectrum of site conditions, including varying equipment standards, operational workflows, and environmental interference, will be necessary to confirm broader applicability. Third, the system assumes a fixed viewpoint in which the interior region of the chute remains clearly visible. If the camera position is unstable, if the chute is occluded by workers, or if the capturing angle significantly deviates from standard placement, the performance of detection and flow computation may degrade. Finally, the dataset primarily reflects construction practices and equipment configurations common in Korea. Further domain adaptation and validation are required to ensure generalizability to regions with different mixer truck designs, operational standards, or construction procedures.